OpenStudy (anonymous):

Solve y=5x^2+25x+3 Show all your work and determine if the graph opens up or down, and if it has a maximum or minimum point.

OpenStudy (cwrw238):

the coefficient of x^2 (5) is positive so it opens upwards and has a minimum value. you can solve this using the quadratic formula.

OpenStudy (cwrw238):

the formula for roots of ax^2 + bx + c = 0 is (-b +/- sqrt(b^2 - 4ac) -------------------- 2a

OpenStudy (anonymous):

Can you please show me the work of you solving this equation?

OpenStudy (anonymous):

Just plug and chug.

OpenStudy (anonymous):

I have no idea what that means please help me

jimthompson5910 (jim_thompson5910):

To solve 5x^2 + 25x + 3 = 0, use the quadratic formula to get \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(25)\pm\sqrt{(25)^2-4(5)(3)}}{2(5)}\] \[\Large x = \frac{-25\pm\sqrt{625-(60)}}{10}\] \[\Large x = \frac{-25\pm\sqrt{565}}{10}\] \[\Large x = \frac{-25+\sqrt{565}}{10} \ \text{or} \ x = \frac{-25-\sqrt{565}}{10}\] \[\Large x \approx -0.123027135199057 \ \text{or} \ x \approx -4.87697286480094\]